# K-theory and localizing invariants of large categories

### v2

@@ 1.12. Colimits of dualizable categories
The second statement follows from the first one by Proposition 1.7.
@@
Let be a directed system of presentable stable categories, and We denote by the transition functors, and by the structural functors to the colimit. Also, denote by and the corresponding right adjoint functors.
1) We have
@@
We use once again the natural equivalence where the transition functors in the inverse system are Moreover, the composition is identified with the functor
To prove 1), we simply observe that for we have isomorphisms
@@
By Proposition 1.66, it is sufficient to prove the statement for
Consider a filtered system of strongly continuous fully faithful functors where the transition functors and are strongly continuous. Put and denote by the colimit of Denote by the structural functors. For we can directly compute the composition Namely, applying Proposition 1.67 together with strong continuity and fully faithfulness of we obtain
@@
Let be a presentable stable category, and suppose we have a directed family of presentable stable full subcategories where the inclusion functors are strongly continuous. If is generated by as a localizing subcategory, then the natural functor is an equivalence.
@@
The following example shows that the analogues of Proposition 1.68 and 1.70 fail in Denote by the -th prime number, i.e. Put and where We consider as a full subcategory of via restriction of scalars, and denote by the inclusion functor. Hence, we have
Note that the subcategories generate as a localizing subcategory. Indeed, we have and However, we claim that the functor is not an equivalence.
Indeed, otherwise by Proposition 1.67 1) we would have But this clearly fails:
@@
The functor commutes with filtered colimits.
@@
The following example shows that Proposition 1.72 does not hold for presentable categories.
@@
Let be a field. Denote by the rectangle (considered as a topological space). Denote by the full subcategory of sheaves whose pullback to vanishes. Then the pullback functors are strongly continuous, the colimit is compactly generated, but we have for
In particular, the functor does not commute with sequential colimits.
@@
It is clear that each functor is strongly continuous. Identifying the colimit with the limit of the inverse sequence, we see that is equivalent to the full subcategory of consisting of sheaves whose support is contained in for all (and whose pullback to vanishes). This is simply the category of sheaves supported at the point which is equivalent to
It remains to show that Since all the categories are equivalent, we may assume Suppose that is a non-zero compact object. For put The pullback functor is strongly continuous, hence is a compact sheaf on Since is contractible and compact, it follows from Proposition 6.14 that is a constant sheaf. Since this holds for arbitrarily small and since we deduce that is a non-zero constant sheaf on . The pullback functor is strongly continuous, hence is also a compact sheaf. But is not compact, a contradiction with Proposition 6.14. ∎
@@
The functor commutes with -filtered colimits. In particular, the functor commutes with -filtered colimits.
@@
Recall that the functor commutes with -filtered colimits. Hence, so is the composition It follows that the functor commutes with -filtered colimits. Passing to compact objects, we deduce that the functor commutes with -filtered colimits.∎
We conclude this subsection with the following elementary statement about pushouts.
@@ 1.13. Semi-orthogonal decompositions
A (finite) semi-orthogonal decomposition in (resp. ) of a presentable stable (resp. dualizable) category is a collection of presentable stable (resp. dualizable) subcategories such that
i) the inclusion functors are strongly continuous;
ii) For we have
iii) The categories generate as a stable subcategory.
In this case we write
@@
If is dualizable, then each is a retract of in hence is dualizable. Conversely, if each is dualizable, then is dualizable by Proposition 1.56.
For the second statement we may assume that Denote by the inclusion functors. Then the functors and satisfy the version of conditions of Remark 1.77 for small stable categories. Hence, they induce a semi-orthogonal decomposition ∎
@@
Note that the functor is automatically strongly continuous.
. It suffices to recall that where is the right adjoint to
. Suppose that is continuous. Note that the functor is given by
@@
Let be a presentable stable category with a semi-orthogonal decomposition in Denote by the left adjoint to and the right adjoint to Let be another presentable stable category.
1) A continuous functor is strongly continuous if and only if the compositions and are strongly continuous. We have a semi-orthogonal decomposition
@@
We first prove part 2). The criterion of strong continuity follows from Proposition 1.9 applied to the functors The functors and are given respectively by precomposition with and The left adjoint to is given by precomposition with and the right adjoint to is given by precomposition with It follows formally that and are fully faithful and the composition is zero. It remains to observe that the cofiber sequence in implies that we have a cofiber sequence in
We now prove part 1). The proof of a semi-orthogonal decomposition is similar; here the inclusion functors are given by the compositions with and The “only if” direction of the criterion is clear. To prove the “if” direction we note that the functors and are strongly continuous, and the cofiber sequence implies that is also strongly continuous. ∎
For a future reference, we define the notion of a poset-indexed (possibly infinite) semi-orthogonal decomposition.
@@
Let be a partially ordered set, and let be a presentable stable (resp. dualizable) category. An -indexed semi-orthogonal decomposition of in (resp. in ) is a collection of presentable stable (resp. dualizable) subcategories such that
i) the inclusion functors are strongly continuous;
ii) For we have
iii) The categories generate as a localizing subcategory.
In this case we write
@@
The first statement is proved in the same way as Proposition 1.78.
To prove the second statement, we introduce for each finite subset the subcategory generated by By Propositions 1.56 and 1.9 the category is dualizable and the inclusion is strongly continuous. Moreover, we have a semi-orthogonal decomposition By Proposition 1.78, the category is generated as a stable category by Combining Corollary 1.70 and Proposition 1.72 we deduce that the category is generated as a stable subcategory by Hence, we have a semi-orthogonal decomposition as required.∎
@@ 1.14. Finite limits of dualizable categories
The limits in the category of dualizable categories are much more complicated then colimits. We start their study with the case of kernels (i.e. fibers).
@@
It is convenient to replace the poset with the isomorphic poset
Choose a sequence in such that and each map is compact. Define the functor by the formula Since any compact morphism in is a composition of two compact morphisms, we can define inductively the functors such that and for any such that the map is compact. The sequence defines a functor
Finally, we define a functor by the formula Then satisfies the required properties (since is a dense linearly ordered set).∎
@@
Let be a dualizable pesentable stable category, and be a full stable subcategory, closed under coproducts. Then there exists the largest dualizable subcategory such that the inclusion functor is strongly continuous.
Moreover, is generated by the objects of the form where is a functor such that for any rational numbers the map is compact in
@@
Denote by the full subcategory generated (as a localizing subcategory) by the objects described in the second assertion. We first prove that is dualizable and the inclusion is strongly continuous.
Consider a functor as above. We may and will assume that for each rational the map is an equivalence. Indeed, as in the proof of Proposition 1.83 we may replace by the functor given by
Now, we observe that our assumptions on imply that each object is contained in (because the linearly ordered set is isomorphic to ). It follows that the functor exists and we have Moreover, the inclusion commutes with hence it is strongly continuous.
Now, let be some full dualizable subcategory, such that the inclusion is strongly continuous. We claim that is contained in To see this, take any countably presented object Choose a functor such that each map is compact in (hence in ) and We see that as required.∎
We can give a description of kernels in
@@
Let be a discrete valuation ring with the maximal ideal and with the field of fractions Then Indeed, the usual kernel is given by but the inclusion functor is strongly continuous. The only localizing subcategory of which is not equal to is zero.
@@
Consider a pair of strongly continuous functors between dualizable categories
1) The fiber product exists in and the natural functor is fully faithful and strongly continuous. More precisely, the category is the largest dualizable subcategory such that the functors and are strongly continuous.
2) If moreover is a localization, then we have and the functor is a localization.
@@
1) By Proposition 1.79 the lax fiber product is dualizable. It follows from Proposition 1.80 that where The usual fiber product (taken in ) is simply The rest follows from Proposition 1.85.
2) It is well-known that in a pullback of a localization is a localization. To show that is also the fiber product in it suffices to show that the functor is a localization (assuming that is a localization). This is standard. Explicitly, the right adjoint is given by and ∎
@@
By Proposition 1.88 we only need to show that the category is compactly generated. It suffices to show that the objects of the form are generated by colimits by the essential image of the diagonal functor
By Proposition 1.20, we may assume that is of the form where and is a zero map in Then we have
@@ 1.15. General limits of dualizable categories
We construct the adjunction unit and counit as follows.
Let be a dualizable stable category. Then the adjunction unit on is given by the functor (we know that this functor is strongly continuous).
Let be an -presentable category. Then the adjunction counit is given by the colimit functor Note that this functor is -strongly continuous.
If is dualizable, then the composition is isomorphic to the identity functor. If is -presentable, then the composition
@@
Given a dualizable category and an -presentable category an -strongly continuous functor corresponds via the above adjunction to the composition
@@
Let be an -category, and be a functor, Then the limit of exists, and is given by
@@
such that the inclusion functor is strongly continuous.
Note that for any functor we have a strongly continuous fully faithful embedding To conclude, we simply observe that is equivalent to
@@
where the latter functor is fully faithful.
This proves the theorem. ∎
@@ 1.16. Example: fpqc descent
To give a feeling of dualizable inverse limits, we prove the following statement. Here we denote by the category of commutatie rings.
@@
1) The functor is a sheaf in the descendable topology.
2) The functor is a sheaf in the fpqc topology.
Recall that the descendable topology on is given by the pretopology in which the covers are the maps such that is a descendable -algebra, see [Mat16, Definition 3.18]. Namely, this means that the unit object is contained in the stable idempotent-complete ideal of generated by
We need the following lemma.
@@
Let be a strongly functor between presentable stable categories, and denote by its right adjoint. Suppose that there exists such that each object of is a direct summand of an object with a finite filtration with at most non-zero subquotients, which are contained in Then for any regular cardinal the functor preserves and reflects -compact objects.
@@
Since is continuous, takes to
If is an object such that is -compact, then for any collection of objects we have
@@
1) Let be a map such that is descendable over Below all the tensor products over are always derived, including the tensor powers of i.e. they are taken in By [Mat16, Proposition 3.22] we have an equivalence By Lemma 1.94, it induces an equivalence By Theorem 1.92, we obtain
It suffices to prove the following statement:
the functor is fully faithful.
To prove recall that by Proposition 1.65 for the complex of morphisms between and in the Calkin category is given by
@@
is an isomorphism. This is equivalent to proving the following
the map is an isomorphism.
This follows easily from the descendability of over Indeed, by [Mat16, Proposition 3.20] the inverse system is pro-constant, hence we have equivalences
2) Descent for finite disjoint unions is obvious. We need to prove that for a faithfully flat map we have As in the proof of 1), we only need to show that and to prove the fully faithfulness statement above.
By Proposition 1.95 below, is isomorphic to an -directed colimit where are descendable faithfully flat -algebras (and is an -directed poset). We know that the functors and commute with -filtered colimits (by Proposition 1.74). Also, the functor commutes with -filtered colimits, since the category is countable. We deduce from part 1) that and the functor is fully faithful. As in the proof of 1), this implies that ∎
@@
Let be a commutative ring. The category of faithfully flat -algebras is -accessible. A faithfully flat -algebra is -compact if and only if is a countably presented -module (equivalently, a countably presented -algebra).
In particular, any faithfully flat -algebra is an -directed colimit of descendable faithfully flat -algebras.
@@
The second assertion follows from the fact that a faithfully flat countably presented -algebra is descendable by [Mat16, Corollary 3.33].
By Lazard’s theorem, the category of flat -modules is equivalent to the ind-completion where is the category of finite projective -modules. The category of flat commutative -algebras is equivalent to the category of modules over the monad The functor commutes with filtered colimits and preserves -compact objects. Applying Proposition 1.98 below, we see that the category is -accessible, and a flat -algebra is -compact if and only if is a countably presented -module.
It remains to observe that if is a map of flat -algebras such that is faithfully flat over then is also faithfully flat over ∎
Remark 1.96.
It seems plausible that for any descendable map and for any uncountable regular we have an equivalence
@@
Let be the category of pairs where Clearly, has filtered colimits and the forgetful functor commutes with filtered colimits. By [Pos23, Theorem 4.1] the category is -accessible, and preserves and reflects -compactness.
Consider as a natural transformation Then the category is equivalent to the joint equifier of the pairs and Applying [Pos23, Theorem 3.1], we conclude that the category is -accessible, and the forgetful functor to preserves and reflects -compactness. ∎

### v3

@@ 1.12. Colimits of dualizable categories
Let be a directed system of presentable stable categories, and We denote by the transition functors, and by the structural functors to the colimit. Also, denote by and the corresponding right adjoint functors. 1) We have
@@
We use once again the natural equivalence where the transition functors in the inverse system are Moreover, the composition is identified with the functor To prove 1), we simply observe that for we have isomorphisms
We prove 2). By Proposition 1.7 the functors are continuous. Using again the continuity of we obtain a well-defined functor given by
Then (1.8) is equivalent to the statement that To check this, take some objects We obtain the following isomorphisms:
Hence, as required.∎
@@
By Proposition 1.65, it is sufficient to prove the statement for Consider a filtered system of strongly continuous fully faithful functors where the transition functors and are strongly continuous. Put and denote by the colimit of Denote by the structural functors. For we can directly compute the composition Namely, applying Proposition 1.66 together with strong continuity and fully faithfulness of we obtain
@@
Let be a presentable stable category, and suppose that we have a directed family of presentable stable full subcategories where the inclusion functors are strongly continuous. If is generated by as a localizing subcategory, then the natural functor is an equivalence.
@@
The following example shows that the analogues of Proposition 1.67 and 1.69 fail in Denote by the -th prime number, i.e. Put and where We consider as a full subcategory of via restriction of scalars, and denote by the inclusion functor. Hence, we have
Note that the subcategories generate as a localizing subcategory. Indeed, we have and However, we claim that the functor is not an equivalence. Indeed, otherwise by Proposition 1.66 1) we would have But this clearly fails:
@@
The functor commutes with filtered colimits. Equivalently, the object is compact.
@@
We now give examples showing that the object is not small (i.e. not -compact for any regular cardinal ). A different argument for this was given previously by German Stefanich, using the generalizations of the derived category of quasi-coherent sheaves on We use the following notation. For a pair of ordinals we denote by the version of a long closed interval: we put
and consider this set as a topological space with the interval topology with respect to the following total order. The order on is lexicographic, and the element is the largest one. Then is a connected compact Hausdorff space.
@@
Let be a field. Take some regular cardinal which we also consider as an ordinal. For an ordinal denote by the rectangle (considered as a topological space). Denote by the full subcategory of sheaves whose pullback to vanishes. Then the pullback functors are strongly continuous for the colimit is non-zero and compactly generated, but we have for Therefore, the functor does not commute with -filtered colimits, i.e. the object is not -compact. In particular, the category is not presentable.
@@
The description of the category implies that it is -presentable: it is a kernel of a continuous functor between dualizable categories with an -continuous right adjoint. It is clear that each functor is strongly continuous. Identifying the colimit with the limit of the inverse system, we see that is equivalent to the full subcategory of consisting of sheaves whose support is contained in for all (and whose pullback to vanishes). This is simply the category of sheaves supported at the point which is equivalent to It remains to show that Since all the categories are equivalent, we may assume Suppose that is a non-zero compact object. For put The pullback functor is strongly continuous, hence is a compact sheaf on Since is connected and compact, it follows from Proposition 6.15 that is a nowhere zero locally constant sheaf. Since this holds for arbitrary we deduce that is a nowhere zero locally constant sheaf on . The pullback functor is strongly continuous, hence is also a compact sheaf. But is not compact, a contradiction with Proposition 6.15. ∎
@@
Let be an uncountable regular cardinal. The functor commutes with -filtered colimits. In particular, the functor commutes with -filtered colimits.
@@
Recall that the functor commutes with -filtered colimits. Hence, so does the composition It follows that the functor commutes with -filtered colimits. Passing to compact objects, we deduce that the functor commutes with -filtered colimits.∎
We conclude this subsection with the following statement about pushouts.
@@ 1.13. Semi-orthogonal decompositions
A (finite) semi-orthogonal decomposition in (resp. ) of a presentable stable (resp. dualizable) category is a collection of presentable stable (resp. dualizable) subcategories such that i) the inclusion functors are strongly continuous; ii) For we have iii) The categories generate as a stable subcategory. In this case we write
@@
If is dualizable, then each is a retract of in hence is dualizable. Conversely, if each is dualizable, then is dualizable by Proposition 1.55. For the second statement we may assume that Denote by the inclusion functors. Then the functors and satisfy the version of conditions of Remark 1.76 for small stable categories. Hence, they induce a semi-orthogonal decomposition ∎
@@
Note that the functor is automatically strongly continuous. . It suffices to recall that where is the right adjoint to . Suppose that is continuous. Note that the functor is given by
@@
Let be a presentable stable category with a semi-orthogonal decomposition in Denote by the left adjoint to and the right adjoint to Let be another presentable stable category. 1) A continuous functor is strongly continuous if and only if the compositions and are strongly continuous. We have a semi-orthogonal decomposition
@@
We first prove part 2). The criterion of strong continuity follows from Proposition 1.8 applied to the functors The functors and are given respectively by precomposition with and The left adjoint to is given by precomposition with and the right adjoint to is given by precomposition with It follows formally that and are fully faithful and the composition is zero. It remains to observe that the cofiber sequence in implies that we have a cofiber sequence in We now prove part 1). The proof of a semi-orthogonal decomposition is similar; here the inclusion functors are given by the compositions with and The “only if” direction of the criterion is clear. To prove the “if” direction we note that the functors and are strongly continuous, and the cofiber sequence implies that is also strongly continuous. ∎
We also define the notion of a poset-indexed (possibly infinite) semi-orthogonal decomposition.
@@
Let be a partially ordered set, and let be a presentable stable (resp. dualizable) category. An -indexed semi-orthogonal decomposition of in (resp. in ) is a collection of presentable stable (resp. dualizable) subcategories such that i) the inclusion functors are strongly continuous; ii) For we have iii) The categories generate as a localizing subcategory. In this case we write
@@
The first statement is proved in the same way as Proposition 1.77. To prove the second statement, we introduce for each finite subset the subcategory generated by By Propositions 1.55 and 1.8 the category is dualizable and the inclusion is strongly continuous. Moreover, we have a semi-orthogonal decomposition By Proposition 1.77, the category is generated as a stable category by Combining Corollary 1.69 and Proposition 1.71 we deduce that the category is generated as a stable subcategory by Hence, we have a semi-orthogonal decomposition as required.∎
@@ 1.14. Finite limits of dualizable categories
The limits in the category of dualizable categories are much more complicated then colimits. We start their study with the case of kernels (i.e. fibers). We denote by the poset of rational numbers with the usual order.
@@
It is convenient to replace the poset with the isomorphic poset Choose a sequence in such that and each map is compact. Define the functor by the formula Since any compact morphism in is a composition of two compact morphisms, we can define inductively the functors such that and for any such that the map is compact. The sequence defines a functor Finally, we define a functor by the formula Then satisfies the required properties (since is a dense linearly ordered set).∎
@@
Let be a dualizable pesentable stable category, and be a full stable subcategory, closed under coproducts. Then there exists the largest dualizable subcategory such that the inclusion functor is strongly continuous. Moreover, is generated by the objects of the form where is a functor such that for any rational numbers the map is compact in
@@
Denote by the full subcategory generated (as a localizing subcategory) by the objects described in the second assertion. We first prove that is dualizable and the inclusion is strongly continuous. Consider a functor as above. We may and will assume that for each rational the map is an equivalence. Indeed, as in the proof of Proposition 1.82 we may replace by the functor given by Now, we observe that our assumptions on imply that each object is contained in (because the linearly ordered set is isomorphic to ). It follows that the functor exists and we have Moreover, the inclusion commutes with hence it is strongly continuous. Now, let be some full dualizable subcategory, such that the inclusion is strongly continuous. We claim that is contained in To see this, take any countably presented object Choose a functor such that each map is compact in (hence in ) and We see that as required.∎
We can now give a description of kernels in
@@
Let be a discrete valuation ring with the maximal ideal and with the field of fractions Then Indeed, the usual kernel is given by but the inclusion functor is not strongly continuous. The only localizing subcategory of which is not equal to is zero.
@@
Consider a pair of strongly continuous functors between dualizable categories 1) The fiber product exists in and the natural functor is fully faithful and strongly continuous. More precisely, the category is the largest dualizable subcategory such that the functors and are strongly continuous. 2) If moreover is a localization, then we have and the functor is a localization.
@@
1) By Proposition 1.78 the lax fiber product is dualizable. It follows from Proposition 1.79 that where The usual fiber product (taken in ) is simply The rest follows from Proposition 1.84. 2) It is well-known that in a pullback of a localization is a localization. To show that is also the fiber product in it suffices to show that the functor is a localization (assuming that is a localization). This is standard. Explicitly, the right adjoint is given by and ∎
@@
By Proposition 1.87 we only need to show that the category is compactly generated. It suffices to show that the objects of the form are generated by colimits by the essential image of the diagonal functor By Proposition 1.19, we may assume that is of the form where and is a zero map in Then we have
@@ 1.15. General limits of dualizable categories
We construct the adjunction unit and counit as follows. Let be a dualizable stable category. Then the adjunction unit on is given by the functor (we know that this functor is strongly continuous). Let be an -presentable category. Then the adjunction counit is given by the colimit functor Note that this functor is -strongly continuous. If is dualizable, then the composition is isomorphic to the identity functor. If is -presentable, then the composition
@@
Given a dualizable category and a -presentable category (where is uncountable), a -strongly continuous functor corresponds via the above adjunction to the composition
@@
Let be an -category, and be a functor, Then the limit of exists, and for any uncountable regular cardinal we have an equivalence
@@
such that the inclusion functor is strongly continuous. Note that for any functor we have a strongly continuous fully faithful embedding To conclude, we simply observe that is equivalent to
@@
where the latter functor is fully faithful. This proves the theorem. ∎
@@ 1.16. Example: fpqc descent
To give a feeling of dualizable inverse limits, we prove two statements about descent for functors with values in We denote by the -category of -ring spectra (without connectivity assumptions). Recall that the descendable topology on is given by the pretopology in which the covers are the maps such that is a descendable -algebra, see [Mat16, Definition 3.18]. Namely, this means that the unit object is contained in the stable idempotent-complete ideal of generated by
@@
The functor is a sheaf in the descendable topology.
This can be deduced from [Ram24b, Corollary 4.85], but we would like to give an explicit proof using the description of the limit from Theorem 1.91. We need the following lemma.
@@
Let be a strongly continuous functor between presentable stable categories, and denote by its right adjoint. Suppose that there exists such that each object of is a direct summand of an object with a finite filtration with at most non-zero subquotients, which are contained in Then for any regular cardinal the functor preserves and reflects -compact objects.
@@
Since is continuous, takes to If is an object such that is -compact, then for any collection of objects we have
@@
Let be a map such that is descendable over By [Mat16, Proposition 3.22] we have an equivalence By Lemma 1.93, it induces an equivalence By Theorem 1.91, we obtain
It suffices to prove the following statement: the functor is fully faithful. To prove recall that by Proposition 1.64 for the spectrum of morphisms between and in the Calkin category is given by
@@
is an isomorphism. This is equivalent to proving the following: the map is an isomorphism. This follows easily from the descendability of over Indeed, by [Mat16, Proposition 3.20] the inverse sequence is pro-constant, hence we have equivalences
This proves the proposition. ∎
Remark 1.94.
It seems plausible that for any descendable map and for any uncountable regular cardinal we have an equivalence
Denote by the category of (discrete) commutatie rings.
Proposition 1.95.
The functor is a sheaf in the fpqc topology.
Proof.
Descent for finite disjoint unions is obvious. We need to prove that for a faithfully flat map we have As in the proof of Proposition 1.92, we only need to show that and to prove the fully faithfulness statement above. By Proposition 1.96 below, is isomorphic to an -directed colimit where are descendable faithfully flat -algebras (and is an -directed poset). We know that the functors and commute with -filtered colimits (by Proposition 1.73). Also, the functor
commutes with -filtered colimits, since the category is countable and the category is compactly generated. We deduce from the proof of Proposition 1.92 that and the functor is fully faithful. This implies that as required. ∎
@@
Let be a commutative ring. The category of faithfully flat -algebras is -accessible. A faithfully flat -algebra is -compact if and only if is a countably presented -module (equivalently, a countably presented -algebra). In particular, any faithfully flat -algebra is an -directed colimit of descendable faithfully flat -algebras.
@@
The last assertion follows from the fact that a faithfully flat countably presented -algebra is descendable by [Mat16, Corollary 3.33]. By Lazard’s theorem, the category of flat -modules is equivalent to the ind-completion where is the category of finite projective -modules. The category of flat commutative -algebras is equivalent to the category of modules over the monad The functor commutes with filtered colimits and preserves -compact objects. Applying Proposition 1.98 below, we see that the category is -accessible, and a flat -algebra is -compact if and only if is a countably presented -module. It remains to observe that if is a map of flat -algebras such that is faithfully flat over then is also faithfully flat over ∎
@@
Let be the category of pairs where Clearly, has filtered colimits and the forgetful functor commutes with filtered colimits. By [Pos23, Theorem 4.1] the category is -accessible, and preserves and reflects -compactness. Consider as a natural transformation Then the full subcategory is exactly the joint equifier of the pairs and Applying [Pos23, Theorem 3.1], we conclude that the category is -accessible, and the forgetful functor to preserves and reflects -compactness. ∎